an unbanked curve has a radius of 60m. coefficient of friction between tyre of tr and road is 0.75. the distance between two front tyres of truck is 2m. if truck exceeds the speed of safe limit then their
  1. inner wheel leaves the ground first
  2. outer wheel leaves the ground first
  3. bboth leave the ground together

Grade:12th

1 Answers

Eshan
5 years ago
Let the normal reactions on the inner and outer wheels beN_1,N_2respectivley.
Let the height of the truck beh.

A pseudo force of\dfrac{mv^2}{r}acts on its center of mass.

Since the body should remain under rotational equilibrium from the point under the center of mass and touching the ground,

N_1(\dfrac{l}{2})+\dfrac{mv^2}{r}.h=N_2(\dfrac{l}{2})
AlsoN_1+N_2=mg

Solving the above equations gives,

N_1=\dfrac{mg}{2}-\dfrac{mv^2h}{lr} andN_2=\dfrac{mg}{2}+\dfrac{mv^2h}{lr}

Hence when speed limit is crossed,N_1becomes zero and the inner tyre leaves the ground.